40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11463     Entry time: Thu Jul 30 03:19:24 2015
Author: ericq 
Type: Update 
Category: LSC 
Subject: Back towards PRFPMI  

The refreshed ALS didn't look so bad today (elog forthcoming), so I decided to give PRFPMI locking a shot tonight. I was able to hold the PRMI while swinging through resonsance, but transitions to RF signals failed. Demod angles / whitening gains/ etc. etc. all need to be rechecked

Some little things here and there that got cleaned up...

  • The PRM oplev beam was being blocked. Why? I removed the block. Should recheck OLTF/spot size on QPD. 
  • ALS -> CARM, DARM signs changed, maybe because I've used the delayed beat as the RF input on the demod board, whereas I imagine it may have been the LO in the beatbox. No big deal.
  • REFL165 whitening gain and input matrix updated. Should recheck demod angles.
  • PRMI triggering settings weren't being set in the script. It's important to include arm transmission signals, since POP2F signals can momentarily dive when swinging through resonance. 
  • I should revisit phase tracker UGF normalization. I/Q amplitudes are varying quite a bit from lock to lock. 
  • PRC angular feedforward disabled for now; need to remeasure the witness/target data with DC coupled ITM oplevs
  • I think there has been a little bit of MC servo tweaking since our last locks, may need to recheck AO TF / gains. 
ELOG V3.1.3-